# Is the $3S$-$2D$ mixing strong for the charmonia $\psi(4040)$ and $\psi(4160)$?

### v1

@@ III An Analysis of Potential Reasons for the Discrepancy
Our obtained results are notably smaller than those reported in Table 1 by other theoretical studies [4, 5, 6, 9, 10, 11]. As shown in Table 1, we can find that derivation of most larger mixing angles relies on fitting the experimentally measured di-electronic widths of the $\psi(4040)$ and $\psi(4160)$, or their ratio. Consequently, the inferred mixing angle is highly dependent on the measured di-electronic widths of the $\psi(4040)$ and $\psi(4160)$. However, when we check the PDG [12], one finds that the reported values of the di-electronic decay width for the $\psi(4160)$ vary significantly across different experimental collaborations. Before 2008, the experimentally measured di-electronic decay width was approximately 0.8 keV [13, 1]. Based on the ratio (1.04) of the di-electronic widths of the $\psi(4040)$ (0.86 keV) and $\psi(4160)$ (0.83 keV), theoretical predictions yielded a mixing angle of about $35^{\circ}$ [4, 6, 5].
@@
In 2008, by considering the interference effect in the analysis of $R$ values, the BES Collaboration reported a higher mass and a smaller di-electronic width for the $\psi(4160)$, which are different from previous results [14]. In 2010, Mo et al. reanalyzed BES data highlighted that the Ref. [14] reported di-electronic width of the $\psi(4160)$ reported is just one of four possible solutions, with the values ranging from 0.4 keV to 1.1 keV [31]. It is crucial to note that the existence of multiple possible solutions for the di-electronic width introduces significant uncertainties into the extraction of the $3S$-$2D$ mixing angle. In Refs. [9, 11] , the authors estimated the $3S$-$2D$ mixing angle to be around $20^{\circ}$ by using a screening potential model and the measured di-electronic width of $0.48\pm 0.22$ keV. A significant $3S$-$2D$ mixing angle was also obtained [10], derived from the ratio (1.79) of the di-electronic widths of the $\psi(4040)$ (0.86 keV) and $\psi(4160)$ (0.48 keV).
@@
Historically, early measurements of the $\psi(4040)$ and $\psi(4160)$ were primarily based on inclusive processes in $e^{+}e^{-}\to\text{hadrons}$, with limited contributions from exclusive decay channels. Especially, the di-electronic decay width of the $\psi(4160)$ extracted from inclusive data exhibits significant uncertainties. This is largely because the numerous open-charm thresholds in this energy region lead to complex structures in the cross-section behavior. The inclusive measurements do not provide sufficient experimental information for a detailed understanding the structure of the high-mass charmonium states. Furthermore, although two exclusive processes involving the $\psi(4160)$ [32, 33], its di-electronic decay width has yet to be precisely determined. Consequently, methods that depend on measured di-electronic widths for determining the mixing angles between $\psi(3^{3}S_{1})$ and $\psi(2^{3}D_{1})$ may lead to unreliable conclusions. Facing these challenges, a more comprehensive understanding of the internal structure and mixing dynamics of the $\psi(4040)$ and $\psi(4160)$ is essential.
@@
To better understand the mixing states $\psi^{\prime}_{3S\text{-}2D}$ and $\psi^{\prime\prime}_{3S\text{-}2D}$, we investigate their OZI-allowed two-body strong decay widths and di-electronic widths, utilizing our calculated mixing angles. The OZI-allowed two-body strong decay width for the $\psi_{3S-2D}^{\prime}$ and $\psi_{3S-2D}^{\prime\prime}$ take the form
@@
Based on the mixing angle, the two-body strong decay decay width of $\psi^{\prime}_{3S\text{-}2D}$ is estimated to be $38.0$ MeV. This value is smaller than the PDG value [12]. However, it is crucial to note that the PDG width for the $\psi(4040)$ is predominantly base on inclusive measurements, rather than exclusive decay channels. Therefore, a direct comparison between our calculated two-body decay width and the PDG’s total width is not straightforward, as the latter encompasses contributions from numerous unmeasured or complex decay modes.
The two-body strong decay width of $\psi^{\prime\prime}_{3S\text{-}2D}$ is predicted to be $31.1$ MeV, which is consistent with the BESIII exclusive measured result $55\pm 15\pm 53$ MeV [33]. Due to the relatively small mixing angle obtained in our model, the dominant decay channel for these states is predicted to remain largely unchanged from that of their unmixed states, supporting the standpoint that their primary two-body strong decay properties are preserved despite mixing.
@@
The di-electronic widths of the mixed $\psi^{\prime}_{3S\text{-}2D}$ and $\psi^{\prime\prime}_{3S\text{-}2D}$, including radiative QCD corrections [34, 35], are given by
@@
Figure 1: The ratio of di-electronic widths $\Gamma^{\psi^{\prime}_{3S\text{-}2D}}_{e^{+}e^{-}}/\Gamma^{\psi^{\prime\prime}_{3S\text{-}2D}}_{e^{+}e^{-}}$ as a function of the $3S\text{-}2D$ mixing angles with our parameters from Eqs. (18)-(19). The red pentagram represents the prediction of this work ($\theta\approx 10^{\circ}$ and ratio $\Gamma^{\psi^{\prime}_{3S\text{-}2D}}_{e^{+}e^{-}}/\Gamma^{\psi^{\prime\prime}_{3S\text{-}2D}}_{e^{+}e^{-}}\approx 4.1$). The light green and dark green shaded bands indicate the experimental values $1.79\pm 0.83$ and $1.04\pm 0.12$, respectively, from the PDG [12]. The triangle and square indicate the mixing angles $20^{\circ}$ and $34.8^{\circ}$, as obtained in Refs. [9] and [6], respectively.
[…]
@@ IV Summary
Despite five decades of research since the $J/\psi$ discovery, the properties of charmonium states above the $D\bar{D}$ threshold, including the $\psi(4040)$ and $\psi(4160)$, remain enigmatic due to experimental and theoretical limitations. While these states are well-documented in the PDG, their underlying physics is not fully understood. An intriguing feature is $S$-$D$ mixing, a phenomenon observed across charmonium states, with former studies proposing $3S$-$2D$ mixing for the $\psi(4040)$ and $\psi(4160)$ [4, 5, 6, 7, 8, 9, 10, 11]. Current determinations of the mixing angle [4, 5, 6, 9, 10, 11], derived from di-electronic widths or their ratios at different experimental periods [13, 14], vary significantly, indicating substantial mixing effects.
This work investigates the dynamical origins of this mixing. Traditional potential models, relying on tensor forces, cannot account for the large observed angles, prompting exploration of alternative mechanisms. Incorporating unquenched effects, particularly coupled-channel dynamics, we calculate mixing angles of $\theta_{1}=7^{\circ}$ and $\theta_{2}=10^{\circ}$ for $\psi^{\prime}_{3S\text{-}2D}$ and $\psi^{\prime\prime}_{3S\text{-}2D}$, respectively—markedly smaller than prior estimates. The discrepancy highlights a critical dependence on experimental di-electronic widths, where current data are inadequate for precise angle extraction. Our results emphasize the necessity of high-precision measurements to resolve the $S$-$D$ mixing puzzle in $\psi(4040)$ and $\psi(4160)$, advancing the broader understanding of charmonium spectroscopy.
@@ Acknowledgements.
X.L. would like to thank Cheng-Ping Shen for his questions following the plenary presentation at the 10th $XYZ$ Symposium held in Changsha, China. This study addresses his questions. This work is supported by the National Natural Science Foundation of China under Grant Nos. 12335001, 12247101, and 12405098, the ‘111 Center’ under Grant No. B20063, the Natural Science Foundation of Gansu Province (No. 22JR5RA389, No. 25JRRA799), the Talent Scientific Fund of Lanzhou University, the fundamental Research Funds for the Central Universities (No. lzujbky-2023-stlt01), and the project for top-notch innovative talents of Gansu province.

### v3

@@ III An Analysis of Potential Reasons for the Discrepancy
Our obtained results are notably smaller than those reported in Table 1 by other theoretical studies [4, 5, 6, 9, 10, 11]. As shown in Table 1, we can find that derivation of most larger mixing angles relies on fitting the experimentally measured dielectronic widths of the $\psi(4040)$ and $\psi(4160)$ or their ratio. Consequently, the inferred mixing angle is highly dependent on the measured dielectronic widths of the $\psi(4040)$ and $\psi(4160)$. However, when we check the PDG [12], one finds that the reported values of the dielectronic decay width for the $\psi(4160)$ vary significantly across different experimental collaborations. Before 2008, the experimentally measured dielectronic decay width was approximately 0.8 keV [13, 1]. Based on the ratio (1.04) of the dielectronic widths of the $\psi(4040)$ (0.86 keV) and $\psi(4160)$ (0.83 keV), theoretical predictions yielded a mixing angle of about $35^{\circ}$ [4, 6, 5].
@@
In 2008, by considering the interference effect in the analysis of $R$ values, the BES Collaboration reported a higher mass and a smaller di-electronic width for the $\psi(4160)$, which are different from previous results [14]. In 2010, Mo, Yuan, and Wang reanalyzed BES data highlighted that the Ref. [14] reported dielectronic width of the $\psi(4160)$ is just one of four possible solutions, with the values ranging from 0.4 to 1.1 keV [31]. It is crucial to note that the existence of multiple possible solutions for the dielectronic width introduces significant uncertainties into the extraction of the 3S3S-$2D$ mixing angle. In Refs. [9, 11] , the authors estimated the 3S3S-$2D$ mixing angle to be around $20^{\circ}$ by using a screening potential model and the measured di-electronic width of $0.48\pm 0.22$ keV. A significant 3S3S-$2D$ mixing angle was also obtained [10], derived from the ratio (1.79) of the dielectronic widths of the $\psi(4040)$ (0.86 keV) and $\psi(4160)$ (0.48 keV).
@@
Historically, early measurements of the $\psi(4040)$ and $\psi(4160)$ were primarily based on inclusive processes in $e^{+}e^{-}\to\text{hadrons}$, with limited contributions from exclusive decay channels. Especially, the dielectronic decay width of the $\psi(4160)$ extracted from inclusive data exhibits significant uncertainties. This is largely because the numerous open-charm thresholds in this energy region lead to complex structures in the cross-section behavior. The inclusive measurements do not provide sufficient experimental information for a detailed understanding the structure of the high-mass charmonium states. Furthermore, although two exclusive processes involving the $\psi(4160)$ [32, 33], its dielectronic decay width has yet to be precisely determined. Consequently, methods that depend on measured dielectronic widths for determining the mixing angles between $\psi(3^{3}S_{1})$ and $\psi(2^{3}D_{1})$ may lead to unreliable conclusions. Facing these challenges, a more comprehensive understanding of the internal structure and mixing dynamics of the $\psi(4040)$ and $\psi(4160)$ is essential.
@@
To better understand the mixing states $\psi^{\prime}_{3S\text{-}2D}$ and $\psi^{\prime\prime}_{3S\text{-}2D}$, we investigate their OZI-allowed two-body strong decay widths and dielectronic widths, utilizing our calculated mixing angles. The OZI-allowed two-body strong decay width for the $\psi_{3S-2D}^{\prime}$ and $\psi_{3S-2D}^{\prime\prime}$ take the form
@@
From Table LABEL:mix-decay, based on the mixing angle, the two-body strong decay decay width of $\psi^{\prime}_{3S\text{-}2D}$ is estimated to be $38.0$ MeV. This value is smaller than the PDG value [12]. However, it is crucial to note that the PDG width for the $\psi(4040)$ is predominantly based on inclusive measurements, rather than exclusive decay channels. Therefore, a direct comparison between our calculated two-body decay width and the PDG’s total width is not straightforward, as the latter encompasses contributions from numerous unmeasured or complex decay modes.
The two-body strong decay width of $\psi^{\prime\prime}_{3S\text{-}2D}$ is predicted to be $31.1$ MeV, which is consistent with the BESIII exclusive measured result $55\pm 15\pm 53$ MeV [33]. Because of the relatively small mixing angle obtained in our model, the dominant decay channel for these states is predicted to remain largely unchanged from that of their unmixed states, supporting the standpoint that their primary two-body strong decay properties are preserved despite mixing.
@@
The dielectronic widths of the mixed $\psi^{\prime}_{3S\text{-}2D}$ and $\psi^{\prime\prime}_{3S\text{-}2D}$, including radiative QCD corrections [34, 35], are given by
@@
Figure 1: The ratio of dielectronic widths $\Gamma^{\psi^{\prime}_{3S\text{-}2D}}_{e^{+}e^{-}}/\Gamma^{\psi^{\prime\prime}_{3S\text{-}2D}}_{e^{+}e^{-}}$ as a function of the $3S\text{-}2D$ mixing angles with our parameters from Eqs. (18)and (19). The red pentagram represents the prediction of this work ($\theta\approx 10^{\circ}$ and ratio $\Gamma^{\psi^{\prime}_{3S\text{-}2D}}_{e^{+}e^{-}}/\Gamma^{\psi^{\prime\prime}_{3S\text{-}2D}}_{e^{+}e^{-}}\approx 4.1$). The light green and dark green shaded bands indicate the experimental values $1.79\pm 0.83$ and $1.04\pm 0.12$, respectively, from the PDG [12]. The triangle and square indicate the mixing angles $20^{\circ}$ and $34.8^{\circ}$, as obtained in Refs. [9] and [6], respectively.
Using a mixing angle of $\theta_{1}=7^{\circ}$, the dielectronic decay width of $\psi^{\prime}_{3S\text{-}2D}$ is calculated to be $0.92~\text{keV}$. This result shows agreement with both the PDG value of $0.86\pm 0.07~\text{keV}$ [12] and the experimentally measured range of $0.6$-$1.4~\text{keV}$ reported in Ref. [31].
For $\theta_{2}=10^{\circ}$, we predict a dielectronic width of $0.22~\text{keV}$ for $\psi^{\prime\prime}_{3S\text{-}2D}$, which approaches the lower experimental bound for the $\psi(4160)$ dielectronic decay width.
@@ IV Summary
Despite five decades of research since the $J/\psi$ discovery, the properties of charmonium states above the $D\bar{D}$ threshold, including the $\psi(4040)$ and $\psi(4160)$, remain enigmatic due to experimental and theoretical limitations. While these states are well documented in the PDG, their underlying physics is not fully understood. An intriguing feature is $S$-$D$ mixing, a phenomenon observed across charmonium states, with former studies proposing 3S3S-$2D$ mixing for the $\psi(4040)$ and $\psi(4160)$ [4, 5, 6, 7, 8, 9, 10, 11]. Current determinations of the mixing angle [4, 5, 6, 9, 10, 11], derived from dielectronic widths or their ratios at different experimental periods [13, 14], vary significantly, indicating substantial mixing effects.
This work investigates the dynamical origins of this mixing. Traditional potential models, relying on tensor forces, cannot account for the large observed angles, prompting exploration of alternative mechanisms. Incorporating unquenched effects, particularly coupled-channel dynamics, we calculate mixing angles of $\theta_{1}=7^{\circ}$ and $\theta_{2}=10^{\circ}$ for $\psi^{\prime}_{3S\text{-}2D}$ and $\psi^{\prime\prime}_{3S\text{-}2D}$, respectively-markedly smaller than prior estimates. The discrepancy highlights a critical dependence on experimental dielectronic widths, where current data are inadequate for precise angle extraction. Our results emphasize the necessity of high-precision measurements to resolve the $S$-$D$ mixing puzzle in $\psi(4040)$ and $\psi(4160)$, advancing the broader understanding of charmonium spectroscopy.
@@ Acknowledgements.
X.L. thanks Cheng-Ping Shen for his questions on my plenary talk at the 10th $XYZ$ Symposium held in Changsha, China. This study addresses his questions. This work is supported by the National Natural Science Foundation of China under Grant No. 12335001, No. 12247101, and No. 12405098, the ”111 Center” under Grant No. B20063, the Natural Science Foundation of Gansu Province (No. 22JR5RA389 and No. 25JRRA799), the Talent Scientific Fund of Lanzhou University, the fundamental Research Funds for the Central Universities (No. lzujbky-2023-stlt01), the project for top-notch innovative talents of Gansu province, and Lanzhou City High-Level Talent Funding.